22.4UC2 Numerical Computation and Statistics in Engineering

Study guide

Classes

Numerical Computation: Thursdays, 4:15 – 5:15, EC 3.03.
Statistics:  Mondays, 4:15 – 5:15, EC 3.03, and

Tuesdays, 2:15 – 3:15, EC 3.03.

Tutorials:  Tuesdays, 10:15 – 11:15, EC 3.01.

Lecturers

Name: Carsten Butz (Statistics).
Office: EC 3.15.
Name: Changhai Wang (Numerical Computations).
Office: EC 2.09.

Textbooks

The following textbooks are recommended. For the statistics part any other good book will do. R. Burden and D. Faires. Numerical Methods. Brookes Cole, 6th edition, 1997. D. G. Rees. Essential Statistics. Chapman & Hall, 1995.

Assingments

There will be three assignments to be handed in during the tutorials. Two of the assignments will cover statistics, the third will cover numerical computations. The assignments are due during the tutorials as follows:

Assignment 1: Week 4,
Assignment 2: Week 6,
Assignment 3: Week 8.

Assessment and Final Examination

The module is assessed through continuous assessment, worth 30% of the total mark, and a written exam, worth 70% of the total mark. The Final Exam will be at the beginning of Term 3. The final mark for this course is the weighted mark of the marks for Assignments and the Final Exam, with weights

each Assignment: 10%,
Final Exam:  70%.

Course Outline

The lecturers will distribute detailed course outlines for their respective parts. Roughly one third of the material will be about Numerical Computations, two thirds about Statistics. This is reflected in the assignments, and will be reflected in the final exam.